The growth of bacteria in a particular food product can be modeled using a Malthusian model with a growth rate of 10% per day. The product cannot be eaten once the bacterial count reaches 3,200,000. If there are initially 400 bacteria present, about how many days will it take before the product is inedible? Setup an equation and solve. Label your answer and round it to the nearest full day.
Answer by stanbon(75887)

The growth of bacteria in a particular food product can be modeled using a Malthusian model with a growth rate of 10% per day. The product cannot be eaten once the bacterial count reaches 3,200,000. If there are initially 400 bacteria present, about how many days will it take before the product is inedible? Setup an equation and solve. Label your answer and round it to the nearest full day.
-------------
3200,000 = 400*(1.10)^x
----
8000 = 1.10^x
----
x = log(8000)/log(1.10) = 94 days when rounded down
----------------
Cheers,
Stan H.
--------------